# X-ray Polarization Detection of the Pulsar Wind Nebula in G21.50.9 with IXPE

### v1

@@ 4 Discussion
The model assuming the same outer polarization structure as seen in radio—where the magnetic field is predominantly radial with respect to the pulsar—predicts that in an annulus with inner and outer radii of ″ and ″, respectively, the PA should be tangentially directed, implying an inferred radial magnetic field, with a background-subtracted and leakage-corrected PD . However, our IXPE analysis finds no significant polarization in this region, strongly suggesting that the magnetic field structures traced by the radio- and X-ray-emitting electrons are different. In this respect, G21.50.9 appears more similar to the Crab Nebula than to Vela, where the radio and X-ray polarization structures are well aligned (Xie et al., 2022; Liu et al., 2023). Interestingly, both G21.50.9 and the Crab Nebula are believed to be in the free-expansion phase, where the PWN expands into the cold SNR ejecta (Bucciantini et al., 2003), and strong Rayleigh-Taylor instabilities shape the outer regions (Jun, 1998; Bucciantini et al., 2004). Vela, on the other hand, is likely in a later evolutionary phase, where interaction with the SNR reverse shock is modifying the structure and dynamics of the nebula (van der Swaluw et al., 2003).
@@
To evaluate the possible contribution of a highly polarized torus, we simulated two different scenarios using ixpeobssim: (i) a uniformly polarized torus embedded in an unpolarized nebula, and (ii) the same torus within a nebula that follows the polarization structure and degree observed in the radio band at 5 GHz (Lai et al., 2022). In both cases, the polarized torus is modeled as an elliptical region (″″) with uniform polarization, inclined by and centered on the Chandra X-ray intensity peak. We varied the torus PD and PA in the simulations until we achieved consistency with the polarization properties measured in the central region of the PWN (space-integrated PCUBE analysis within a radius ″ from the PSR: PD , PA ). This region was chosen because it provides the highest polarization significance while minimizing leakage effects, which are not simulated by ixpeobssim (see Appendix B.2). The results of these simulations for the two model configurations are shown in Table 3. Both models yield large uncertainties in the intrinsic polarization properties of the torus, and neither scenario is strongly favored, though the unpolarized nebula case appears somewhat closer to the infrared observations. In either case, the results are consistent with previous IXPE studies of other PWNe (Bucciantini et al., 2023b; Romani et al., 2023; Xie et al., 2022), suggesting that even in the case of G21.50.9, close to the termination shock—where particles are likely being accelerated to high energies—turbulence remains moderate ( at most ), indicating a relatively ordered magnetic field.
@@ Appendix A Spectro-polarimetric analysis
In this section of the Appendix, we provide additional plots to support the spectro-polarimetric analysis of the entire PWN, as described in Sect. 3.1. Fig. 4 presents the Stokes I, Q and U spectra for the three DUs, along with their best-fit models and the residuals from the 3ML analysis. While the Stokes Q and U spectra are well described by the model, the Stokes I spectral fit shows significant deviations at low energies, most likely due to instrumental calibration issues. These discrepancies become more noticeable as the statistical uncertainties approach the level of systematic errors. To assess the potential impact of these issues on the polarization measurement, we repeated the same analysis in two additional configurations: (i) applying the off-axis vignetting correction to the on-axis IXPE effective area (version 20240125, used throughout the paper) using the HEASARC tool ixpecalcarf666https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/help/ixpecalcarf.py.html, and (ii) employing the previous time-independent version of the IXPE IRFs (version 20230526, released on June 16, 2023). In both cases, while the best-fit power-law index varied within the range 1.85–1.98, the resulting PD and PA values remained virtually unchanged. Overall, despite the minor calibration uncertainties that affect the spectral analysis, the stability of the polarization measurements confirms the robustness and reliability of our polarimetric results.
@@ Appendix B Polarization leakage evaluation and correction
Fig. 5 shows the uncorrected polarization map of G21.50.9 resulting from the space-resolved analysis (2–8 keV). Compared to Fig. 3, an outflow pattern is clearly visible, caused by the superposition of the intrinsic, nearly uniform polarization of the source and the radial footprint due to leakage. The purpose of this appendix is to review and compare the various methods used to evaluate or subtract the effect of the polarization leakage. As reported in Sect. 3.2, we employed four different techniques: the LeakageLib software package, a full ixpesim/ixpeobssim simulation of the source, a joint machine learning and analytic event reconstruction method, and the Mueller matrix approach, each one described in the following sections. Although each method has some limitations, LeakageLib is currently the most advanced correction technique. For this reason, we used it to generate Fig. 3 and adopted it as the reference for comparison with the other methods (Fig. 6). Overall, all four methods provide consistent results, reinforcing confidence in our polarization leakage correction techniques. The LeakageLib and Mueller matrix methods yield nearly identical outcomes, while the hybrid event reconstruction provides an independent confirmation by directly mitigating leakage effects. The ixpesim/ixpeobssim simulation slightly overestimates leakage correction but remains within acceptable limits. These comparisons highlight the robustness of our polarization analysis and the importance of accurately modeling leakage effects in extended sources like G21.50.9.
@@ B.1 LeakageLib
For extended sources, the tool provides routines to estimate the expected Stokes maps (I, Q and U) due to polarization leakage based on a Chandra observation of the source. The Chandra count map is assumed to represent the true morphology given its far superior angular resolution compared to IXPE. However, in the current release (v1.1.0), no PSF dependence on off-axis angle or event energy has yet been implemented. For G21.50.9, we modeled the source using the same Chandra map employed throughout the paper. The output Stokes maps, containing the predicted signal induced by leakage, are then normalized to match the observed data and properly rebinned using the same grid as in Fig. 5. For all IXPE telescopes, the content of the Stokes Q and U maps is finally subtracted from the polarization map obtained in Sect. 3.2.
@@
The ixpeobssim (Baldini et al., 2022) software allows the user to simulate an IXPE observation starting from a generic model of the source or a Chandra observation. However, ixpeobssim alone cannot predict polarization leakage, as this effect is not included in the standard IRFs. To enable leakage prediction, ixpeobssim offers an interface to ixpesim (Di Lalla, 2019), which implements a detailed photon-by-photon GEANT4 simulation of the physical interactions in the GPD. As described in the official documentation777https://ixpeobssim.readthedocs.io/en/latest/ixpesim.html, a list of photons can be simulated in ixpeobssim up to the Beryllium window on top of the GPD using the xpphotonlist application. These photons are then propagated throughout the detector using ixpesim and turned into a list of actual photoelectron tracks for the events that trigger the detector. The output file containing the track images (comparable to the IXPE Level-1 data) can be reconstructed using the same event reconstruction algorithm used for the flight data, and a final step using xpsimfmt, part of ixpeobssim, closes the loop by formatting it in a way virtually identical to the actual IXPE Level-2 data.
Based on a Chandra observation of G21.50.9 (Obs ID: 5166), we ran 100 full ixpesim/ixpeobssim simulations, equivalent to a total exposure time of 40 Ms. We modeled the source as intrinsically unpolarized, so that any pattern detected in the output polarization map could be attributed to the leakage effect. By rescaling and rebinning the resulting Stokes maps with the same binning scheme as in Fig. 5, their contribution can be subtracted pixelwise to correct the polarization leakage effect. Fig. 6 shows a detailed comparison between the outcome of LeakageLib and the other three methods. In particular, top row is relative to the ixpesim/ixpeobssim simulation, and from left to right, the three columns show the corrected polarization map obtained with this method (2–8 keV) and the differences in sigmas of the maps of degree and angle of polarization compared to LeakageLib. In contrast to all other methods, this approach appears to over-correct the leakage effect, inducing a spurious tangential pattern. This could be due to an imperfect tuning of the detector parameters in ixpesim, given the observed time evolution of the GPD. Nevertheless, the differences measured between the two methods are within 3, although clearly spatially correlated due to the over-prediction of the leakage amplitude by the simulation.
@@ B.3 Hybrid event reconstruction
The hybrid event reconstruction is a recently developed method, described in detail in Cibrario et al. (2023), which combines a new machine learning-based prediction of the photon impact point with a standard moment analysis reconstruction of the photoelectron track, commonly used for IXPE data. This joint approach exploits the better performance achieved by convolutional neural networks in predicting the conversion points from the track images, improving the IXPE polarization capabilities (expressed by the modulation factor) and partially mitigating the effect of polarization leakage (Cibrario et al. in prep).
@@ B.4 Mueller matrix formalism
The Mueller matrix approach, thoroughly discussed in Bucciantini et al. (2023a), is essentially a generalization of the PSF for polarized observations, where the elements of the Mueller matrix describe how the intrinsic Stokes parameters mix with each other due to the effect of polarization leakage. The primary assumption of this method is that the displacement of the reconstructed absorption point from the true position is correlated with the direction of the reconstructed polarization plane (preferentially displaced along the direction of the polarization plane). This simplifies the problem to a linear displacement model. The matrix elements can be derived either theoretically from a given PSF, or by fitting the expected functional form using in-flight IXPE data. In either case, computing the Stokes maps of any source is reduced to a simple convolution with a Chandra count map. For G21.50.9, we adopted parameters derived from the IXPE observation of Cygnus X-1 (Obs ID: 01002901), one of the brightest, weakly polarized point sources observed by IXPE. Unlike LeakageLib, the tool implementing this formalism assumes a circular, azimuthally symmetric PSF, as the one provided by ixpeobssim. Although this is an approximation of the real PSF (see Dinsmore & Romani (2024) for a detailed comparison), this model still provides a reasonable estimate of the average effect across the three IXPE telescopes. As previously done for LeakageLib, the leakage estimation is based on a Chandra observation of G21.50.9, and the resulting Stokes maps are properly normalized, rebinned and finally subtracted from the uncorrected polarization map, to allow a meaningful comparison. As for the other methods, the subtracted polarization map and its comparison with LeakageLib is reported in Fig. 6 (bottom row). The results of these two methods are in excellent agreement, with differences in the PD and PA maps well below 1, most likely due to differences in PSF parameterization.

### v2

@@ 4 Discussion
The model that assumes the same outer polarization structure observed in the radio band—where the magnetic field is predominantly radial with respect to the pulsar—predicts that, in an annulus with inner and outer radii of ″ and ″, respectively, the PA should be tangential, with a background-subtracted and leakage-corrected PD . Using the xpstokesalign tool of ixpeobssim, we aligned the reconstructed Stokes parameters of this region to a tangential (radio-like) polarization model by rotating the Stokes vectors event by event, based on their spatial position. Despite an MDP of , we did not find evidence of significant polarization. This result strongly suggests that the magnetic field structures traced by the radio- and X-ray-emitting electrons are different. In this respect, G21.50.9 appears more similar to the Crab Nebula than to Vela, where the radio and X-ray polarization structures are well aligned (Xie et al., 2022; Liu et al., 2023). Interestingly, both G21.50.9 and the Crab Nebula are believed to be in the free-expansion phase, where the PWN expands into the cold SNR ejecta (Bucciantini et al., 2003), and strong Rayleigh-Taylor instabilities shape the outer regions (Jun, 1998; Bucciantini et al., 2004). Vela, on the other hand, is likely in a later evolutionary phase, where interaction with the SNR reverse shock is modifying the structure and dynamics of the nebula (van der Swaluw et al., 2003).
@@
To evaluate the possible contribution of a highly polarized torus, we simulated two different scenarios using ixpeobssim: (i) a uniformly polarized torus embedded in an unpolarized nebula, and (ii) the same torus within a nebula that follows the polarization structure and degree observed in the radio band at 5 GHz (Lai et al., 2022). In both cases, the polarized torus is modeled as an elliptical region (″″) with uniform polarization, centered on the CXO X-ray intensity peak and oriented with a position angle on the plane of the sky of , consistent with the measured PA and predictions from magneto-hydrodynamic models. We varied the torus PD and PA in the simulations until we achieved consistency with the polarization properties measured in the central region of the PWN (space-integrated PCUBE analysis within a radius ″ from the PSR: PD , PA ). This region was chosen because it provides the highest polarization significance while minimizing leakage effects, which are not simulated by ixpeobssim (see Appendix B.2). The results of these simulations for the two model configurations are shown in Table 3. Both models yield large uncertainties in the intrinsic polarization properties of the torus, and neither scenario is strongly favored, though the unpolarized nebula case appears somewhat closer to the infrared observations. In either case, the results are consistent with previous IXPE studies of other PWNe (Bucciantini et al., 2023b; Romani et al., 2023; Xie et al., 2022), suggesting that even in the case of G21.50.9, close to the termination shock—where particles are likely being accelerated to high energies—turbulence remains moderate ( at most ), indicating a relatively ordered magnetic field.
The PA measured by IXPE for G21.50.9 in the space-integrated analysis differs by approximately from the spin-axis orientation of PSR J18331034 inferred from morphological modeling of the X-ray torus (Ng & Romani, 2008). This discrepancy may be explained by the fact that the IXPE measurement reflects the integrated polarization from the entire PWN, including contributions from the more extended outer regions, rather than isolating the compact torus. As demonstrated by our simulations (Table 3), even small residual polarization from the outer nebula or deviations from perfect axisymmetry can bias the inferred PA away from the spin-axis direction. A similar effect has been observed in the Crab nebula (Bucciantini et al., 2023b), where non-axisymmetric surface brightness in the X-ray torus produces a measurable offset between PA and the PWN symmetry axis. Given these considerations and the modest statistical tension, we regard the measured PA and spin-axis orientation as broadly consistent within uncertainties.
@@ Appendix A Spectro-polarimetric analysis
In this section of the Appendix, we provide additional plots to support the spectro-polarimetric analysis of the entire PWN, as described in Sect. 3.1. Fig. 4 presents the Stokes I, Q and U spectra for the three DUs, along with their best-fit models and the residuals from the 3ML analysis. While the Stokes Q and U spectra are well described by the model, the Stokes I spectral fit shows significant deviations at low energies, most likely due to instrumental calibration issues. These discrepancies become more noticeable as the statistical uncertainties approach the level of systematic errors. To assess the potential impact of these issues on the polarization measurement, we repeated the same analysis in two additional configurations: (i) applying the off-axis vignetting correction to the on-axis IXPE effective area (version 20240125, used throughout the paper) using the HEASARC tool ixpecalcarf666https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/help/ixpecalcarf.py.html, and (ii) employing the previous time-independent version of the IXPE IRFs (version 20230526, released on June 16, 2023). In both cases, while the best-fit power-law index varied within the range 1.85–1.98, the resulting PD and PA values remained virtually unchanged.
The IXPE energy response matrix is defined in 275 channels spanning 1–12 keV (0.04 keV per bin), and we analyzed the resulting spectra without additional re-binning, as in other IXPE studies (Negro et al., 2023). We verified that rebinning the data to coarser energy grids (e.g., 25 or 100 bins over 2–8 keV) does not affect the spectro-polarimetric results. Given the featureless, non-thermal nature of the source spectrum, we find no need to apply optimized rebinning schemes such as that proposed by Kaastra & Bleeker (2016). Overall, despite the minor calibration uncertainties that affect the spectral analysis, the stability of the polarization measurements confirms the robustness and reliability of our polarimetric results.
@@ Appendix B Polarization leakage evaluation and correction
Fig. 5 shows the uncorrected polarization map of G21.50.9 obtained from the space-resolved analysis (2–8 keV). Compared to Fig. 3, the outer regions appear to exhibit an outflow pattern, which results from the superposition of the intrinsic, nearly uniform polarization of the source and the radial footprint introduced by polarization leakage. In this appendix, we provide a detailed comparison of the different techniques used to evaluate and correct the effect of polarization leakage, which is particularly relevant for spatially resolved polarimetric analyses of extended sources such as G21.50.9. The purpose of this comparison is to validate the robustness of the recovered polarization morphology shown in Fig. 3.
As reported in Sect. 3.2, four independent techniques were employed: the LeakageLib software package, a full ixpesim/ixpeobssim simulation of the source, a joint machine learning and analytic event reconstruction method, and the Mueller matrix approach, each one described in the sections below. Although each method has some limitations, LeakageLib is currently the most advanced correction technique and was used to generate Fig. 3. Fig. 6 provides a side-by-side comparison of the polarization maps obtained using the three alternative methods relative to LeakageLib, chosen as a reference. The leftmost column of each row shows the corrected polarization map using the indicated method, while the middle and right columns show the residuals in PD and PA compared to LeakageLib, expressed in units of combined uncertainty. The LeakageLib and Mueller matrix methods yield nearly identical outcomes, while the hybrid event reconstruction provides an independent confirmation by directly mitigating leakage effects. The ixpesim/ixpeobssim simulation slightly overestimates leakage correction but remains within acceptable limits. While small variations exist at the edges of the nebula, where leakage and systematic uncertainties become more prominent, the general structure is stable across methods. Overall, all four techniques yield broadly consistent polarization patterns—especially in the bright, central region of the PWN where the statistical significance is highest—reinforcing confidence in our polarization leakage correction techniques and the robustness of our polarization analysis.
@@ B.1 LeakageLib
For extended sources, the tool provides routines to estimate the expected Stokes maps (I, Q and U) due to polarization leakage based on a CXO observation of the source. The CXO count map is assumed to represent the true morphology given its far superior angular resolution compared to IXPE. However, in the current release (v1.1.0), no dependence on the IXPE PSF off-axis angle or event energy has yet been implemented. For G21.50.9, we modeled the source using the same CXO map employed throughout the paper. The output Stokes maps, containing the predicted signal induced by leakage, are then normalized to match the observed data and properly rebinned using the same grid as in Fig. 5. For all IXPE telescopes, the content of the Stokes Q and U maps is finally subtracted from the polarization map obtained in Sect. 3.2.
@@
The ixpeobssim (Baldini et al., 2022) software allows the user to simulate an IXPE observation starting from a generic model of the source or a CXO observation. However, ixpeobssim alone cannot predict polarization leakage, as this effect is not included in the standard IRFs. To enable leakage prediction, ixpeobssim offers an interface to ixpesim (Di Lalla, 2019), which implements a detailed photon-by-photon GEANT4 simulation of the physical interactions in the GPD. As described in the official documentation777https://ixpeobssim.readthedocs.io/en/latest/ixpesim.html, a list of photons can be simulated in ixpeobssim up to the beryllium window on top of the GPD using the xpphotonlist application. These photons are then propagated throughout the detector using ixpesim and turned into a list of actual photoelectron tracks for the events that trigger the detector. The output file containing the track images (comparable to the IXPE Level-1 data) can be reconstructed using the same event reconstruction algorithm used for the flight data, and a final step using xpsimfmt, part of ixpeobssim, closes the loop by formatting it in a way virtually identical to the actual IXPE Level-2 data.
Based on a CXO observation of G21.50.9 (Obs ID: 5166), we ran 100 full ixpesim/ixpeobssim simulations, equivalent to a total exposure time of 40 Ms. We modeled the source as intrinsically unpolarized, so that any pattern detected in the output polarization map could be attributed to the leakage effect. By rescaling and rebinning the resulting Stokes maps with the same binning scheme as in Fig. 5, their contribution can be subtracted pixelwise to correct the polarization leakage effect. The top row of Fig. 6 shows a detailed comparison between the outcome of the ixpesim/ixpeobssim simulation technique and LeakageLib. In contrast to all other methods, this approach appears to over-correct the leakage effect, inducing a spurious tangential pattern. Previous validation studies involving simpler cases, such as point sources, have shown that the ixpesim/ixpeobssim approach tends to systematically overestimate the level of polarization leakage, likely due to an imperfect tuning of the detector parameters in ixpesim, given the observed time evolution of the GPD. Nevertheless, the differences measured between the two methods are within 3, although clearly spatially correlated due to the over-prediction of the leakage amplitude by the simulation.
@@ B.3 Hybrid event reconstruction
The hybrid event reconstruction is a recently developed method, described in detail in Cibrario et al. (2023), which combines a new machine learning-based prediction of the photon impact point with a standard moment analysis reconstruction of the photoelectron track, commonly used for IXPE data. This joint approach exploits the better performance achieved by convolutional neural networks in predicting the conversion points from the track images, improving the IXPE polarization capabilities (expressed by the modulation factor) and partially mitigating the effect of polarization leakage (Cibrario et al., 2025).
@@ B.4 Mueller matrix formalism
The Mueller matrix approach, thoroughly discussed in Bucciantini et al. (2023a), is essentially a generalization of the PSF for polarized observations, where the elements of the Mueller matrix describe how the intrinsic Stokes parameters mix with each other due to the effect of polarization leakage. The primary assumption of this method is that the displacement of the reconstructed absorption point from the true position is correlated with the direction of the reconstructed polarization plane (preferentially displaced along the direction of the polarization plane). This simplifies the problem to a linear displacement model. The matrix elements can be derived either theoretically from a given PSF, or by fitting the expected functional form using in-flight IXPE data. In either case, computing the Stokes maps of any source is reduced to a simple convolution with a CXO count map. For G21.50.9, we adopted parameters derived from the IXPE observation of Cygnus X-1 (Obs ID: 01002901), one of the brightest, weakly polarized point sources observed by IXPE. Unlike LeakageLib, the tool implementing this formalism assumes a circular, azimuthally symmetric PSF, as the one provided by ixpeobssim. Although this is an approximation of the real PSF (see Dinsmore & Romani (2024) for a detailed comparison), this model still provides a reasonable estimate of the average effect across the three IXPE telescopes. As previously done for LeakageLib, the leakage estimation is based on a CXO observation of G21.50.9, and the resulting Stokes maps are properly normalized, rebinned and finally subtracted from the uncorrected polarization map, to allow a meaningful comparison. As for the other methods, the subtracted polarization map and its comparison with LeakageLib is reported in Fig. 6 (bottom row). The results of these two methods are in excellent agreement, with differences in the PD and PA maps well below 1, most likely due to differences in PSF parameterization.